The line passes through ​(9​,−3​) and is parallel to the line whose equation is y=2x+4.
Answer by addingup(3677)

y = mx+b
y = 2x+4 set x to 0 and solve for y:
(0, 4)
Your slope, m, is 2. Slope is rise/run, in this case rise = 2 and run = 1 because every number is divisible by itself and by 1. In other words:
y = 2/1x+4
What does this mean, rise over run? The rise is up and the run is across, horizontal.
The rise says that our y is 2 above the intercept, and the run is our x:
(1, 2+4) = (1, 6)
Now you have an ordered pair: (0, 4)(1, 6) Graph your line. Then place a dot at (9, -3), and finally find another set of points (there are many, any of them is correct) that would make the line paralel to the line given by (0, 4)(1, 6)